Under which reaction conditions could (delta)H equal (delta)E? (provide atleast two conditions)
Answer #1

H = E +PV
ΔH = Δ(E + PV)
ΔH = ΔE = Δ(PV)

At constant pressure:

ΔH = ΔE + PΔV

So if you are at constant pressure and constant volume, then:

ΔH = ΔE + P(0)
ΔH = ΔE

If we don't have ΔV but we have other info (still at constant P)

PV = nRT

V = nRT/P

If we are at constant P and T then only n will vary giving us:

ΔV = (Δn)(nRT/P)

Therefore:

ΔH = ΔE + PΔV
ΔH = ΔE + P(Δn)(RT/P)
ΔH = ΔE + (Δn)RT

So from here if we are talking about any reaction and the change in no. of moles of the reaction is zero.

A -----> B

one mole of A gives one mole of B, change in no. of moles at constt T , P is zero, in that case also ΔH = ΔE.